The productive capability of an economy is such that to produce 5 units of military good it takes 2 workers to be employed while 10 units of consumer goods require 3 workers. Resources are limited in such a way that only 75 units of military good can be produced when all resources are employed.

a) How much workers are in the economy?

b) What is the maximum amount of consumer goods that can be produced?



1
Expert's answer
2021-05-17T10:38:46-0400

Number of military goods produced per worker could produce

"=\frac{Number\space of\space military\space goods\space produced}{Number\space of\space workers\space taken \space to \space produce\space it }"


"=\frac{5\space units\space of \space military\space \space goods}{2\space workers}"


=2.5 units of military goods per worker


Number of workers used for production

"=\frac{Number\space of\space military \space goods\space produced}{Number\space of\space military\space goods \space produced\space per\space worker}"


"=\frac{75\space units\space of \space goods}{2.5\space military\space goods\space per\space worker}"


= 30 workers


Number of consumer goods produced per worker could produce

"=\frac{Number\space of\space consumer\space goods\space produced}{Number\space of\space workers\space taken\space to \space produce \space it}"


"=\frac{10\space units\space of\space consumer\space goods}{3\space workers}"


= 3.33 units of consumer goods per worker.


Number of consumer goods produced

"=Number\space of\space workers\space used\space for \space production \times Number\space of\space consumer\space goods\space produced\space per\space worker"


"=30\space workers\times 3.33\space consumer\space goods\space per\space worker"

"=99.99\space units\space of\space consumer\space goods\space produced."

"\approx100\space units\space of\space consumer\space goods\space produced"


a.There are 30 workers in the economy.

b. There are 100 units of consumer goods produced in the economy.